# From Synthetic to Real: Unveiling the Power of Synthetic Data for Video Person Re-ID

### v2

@@ Methods
CAViT (Wu et al., 2022a) 13.8 24.9 - 6.5 4.0 7.3 - 6.7
PTSA (Wang et al., 2021) 34.2 50.2 56.4 20.1 23.7 42.3 52.3 30.5
AP3D (Gu et al., 2020) 44.7 61.3 66.5 25.0 22.7 39.3 50.0 31.7
TCLNet (Hou et al., 2020) 42.6 57.7 64.9 22.8 26.7 50.0 61.3 34.9
CAL (Gu et al., 2022) 43.6 58.3 63.6 23.4 25.3 41.3 51.7 31.7
SINet (Bai et al., 2022) 42.8 57.6 63.1 23.4 29.3 47.3 63.3 39.8
GLT (Zheng et al., 2021a) 38.5 55.3 62.9 24.0 12.0 26.0 34.0 20.0
QAconv (Liao and Shao, 2022) 40.7 55.6 62.4 22.7 20.0 38.0 48.0 29.4
DCCT (Liu et al., 2023) 42.8 58.9 65.3 23.3 24.7 45.3 61.6 32.4
S2P (Rami et al., 2024) 32.4 48.9 55.2 19.7 18.7 35.3 43.4 27.0
TF-CLIP (Chenyang Yu, 2024) 45.7 60.5 67.5 25.7 31.3 53.7 68.7 42.1
@@
CAViT (Wu et al., 2022a) 13.5 38.2 - 25.2 12.2 20.5 - 4.1
PTSA (Wang et al., 2021) 47.9 71.4 80.1 51.8 10.6 18.1 22.5 3.7
AP3D (Gu et al., 2020) 56.2 80.9 92.1 67.4 13.4 21.7 26.1 4.5
TCLNet (Hou et al., 2020) 65.6 74.5 77.9 57.5 11.0 19.2 23.6 3.9
CAL (Gu et al., 2022) 51.6 73.0 78.7 55.1 11.4 20.3 25.2 4.0
SINet (Bai et al., 2022) 53.7 75.3 83.2 63.0 13.8 21.7 26.4 4.7
GLT (Zheng et al., 2021a) 33.7 62.9 71.9 48.0 9.1 16.7 20.9 3.1
QAconv (Liao and Shao, 2022) 50.6 77.5 84.3 61.8 11.3 20.6 25.6 4.1
DCCT (Liu et al., 2023) 49.7 74.7 81.4 59.5 14.3 22.5 28.3 5.3
S2P (Rami et al., 2024) 40.4 67.5 84.3 53.5 10.1 18.1 22.3 3.4
TF-CLIP (Chenyang Yu, 2024) 63.4 72.4 88.7 67.8 16.2 25.6 31.8 5.7
@@
Table 2: Comparison with the state-of-the-art Re-ID methods from the synthetic SVReID with clothes changing to the real dataset CCVID. (%)
@@
CAViT (Wu et al., 2022a) 61.4 70.0 - 42.5
PTSA (Wang et al., 2021) 61.4 74.3 80.0 52.6
TCLNet (Hou et al., 2020) 42.8 57.6 63.1 23.4
CAL (Gu et al., 2022) 70.7 78.1 81.6 66.8
SINet (Bai et al., 2022) 70.7 79.1 83.1 64.1
GLT (Zheng et al., 2021a) 68.2 77.6 80.9 59.1
QAconv (Liao and Shao, 2022) 55.4 70.4 75.3 37.8
AP3D (Gu et al., 2020) 52.8 62.6 68.2 43.1
DCCT (Liu et al., 2023) 66.7 78.9 80.3 57.8
S2P (Rami et al., 2024) 53.4 65.7 72.3 41.3
TF-CLIP (Chenyang Yu, 2024) 72.2 81.3 84.5 68.3
@@
Table 3: Ablation study of the proposed method. (%)
@@
+ $\mathcal{L}^{\mathrm{S}}_{1}$ 32.7 59.3 73.3 45.4 71.7 78.3 81.8 68.1
+$\mathcal{L}^{\mathrm{S}}_{1}$ + $\mathcal{L}^{\mathrm{S}}_{2}$ 34.0 62.7 75.3 47.3 72.9 80.5 83.1 68.2
+ $\mathcal{L}^{\mathrm{S}}_{1}$ + $\mathcal{L}^{\mathrm{S}}_{2}$ + $\mathcal{L}^{\mathrm{S}}_{3}$ 36.0 63.3 75.3 49.1 73.9 79.9 84.2 70.1
w/o $\mathcal{L}^{\mathrm{C}}_{1}$ 39.0 62.7 73.3 50.1 75.2 82.4 86.8 71.8
w/o $\mathcal{L}^{\mathrm{C}}_{2}$ 39.3 64.0 78.0 51.2 74.6 80.3 83.8 70.0
@@
4.2 Comparison with state-of-the-art methods
Comparison methods. Due to limited research on cross-domain person Re-ID based on videos, we conducted a comprehensive evaluation by comparing our method against two categories of approaches. The first category is the state-of-the-art video-based person Re-ID approaches, including CAViT (Wu et al., 2022a), PSTA (Wang et al., 2021), AP3D (Gu et al., 2020), TCLNet (Hou et al., 2020), SINet (Bai et al., 2022), CAL (Gu et al., 2022), DCCT (Liu et al., 2023) and TF-CLIP (Chenyang Yu, 2024). The second category is image-based person Re-ID methods that aim to handle the unsupervised domain adaptation (UDA) problem, including GLT (Zheng et al., 2021a), QAconv (Liao and Shao, 2022) and S2P (Rami et al., 2024).
As shown in Table 1, we compare our method with the state-of-the-art video person Re-ID approaches and domain adaptive Re-ID approaches. We conduct the experiments on four pairs of cross-domain datasets. We can first see that the proposed method gets the best performance compared to other approaches with a large margin, on all four real-world datasets. Specifically, we find that our method outperforms the second-best approach by over 10% on iLIDS-VID and PRID datasets. The performance improvement on the large-scale Mars dataset also achieves over 5%. We can also see that on LS-VID the results of all methods are generally poor. This is because LS-VID is a large-scale dataset with more challenging testing cases.
As shown in Table 2, on the CCVID dataset, the clothes-changing setting increases the difficulty for Re-ID. But since the data variety of CCVID (like the view angles, and illumination) is relatively simple, the overall performance on it is generally high. In this case, the proposed method still performs better than CAL, which is specifically designed for CCVID to handle the clothes-changing problem. These results demonstrate that the proposed method has an effective domain adaption ability across synthetic and real-world datasets.
4.3 Ablation study
To verify the effectiveness of each component in the proposed method, we conduct the ablation study on two datasets, i.e., the medium-scale dataset iLIDS-VID and the clothes-changing aware dataset CCVID, as shown in Table 3.
Effectiveness of domain-invariant feature learning module (Section 3.2). We add the self-supervised domain discrimination losses, i.e., $\mathcal{L}^{\mathrm{S}}_{1}$, $\mathcal{L}^{\mathrm{S}}_{2}$, and $\mathcal{L}^{\mathrm{S}}_{3}$ in Eq. (10) one by one. As shown in Table 3, we can first see that the proposed self-supervised loss $\mathcal{L}^{\mathrm{S}}$ is effective in improving the Re-ID performance with a large margin. Also, the results demonstrate that each part in $\mathcal{L}^{\mathrm{S}}$ can improve the performance, respectively.
Table 4: Comparison results with the real-to-real data training. (%)
@@ 4.3 Ablation study
Effectiveness of cross-domain consistency learning module (Section 3.3). With $\mathcal{L}^{\mathrm{S}}$ equipped, we further investigate the effectiveness of the mean-teacher scheme. We can first see that the integration of the mean-teacher training strategy can further improve the performance of our method. Also, we remove each corresponding loss component i.e., $\mathcal{L}^{\mathrm{C}}_{1}$ in Eq (14) and $\mathcal{L}^{\mathrm{C}}_{2}$ in Eq. (15) in the self-supervised ID consistency learning, respectively. We can also see that each sub-loss in $\mathcal{L}^{\mathrm{C}}$ is verified to be useful.
@@
4.4 Cross-dataset evaluation results
In this section, we also use two real datasets, i.e., iLIDS-VID and CCVID for testing, and use the other three real datasets, i.e., Mars, PRID, and LS-VID for training. As shown in Table 4.3, we compare the cross-domain Re-ID results using the real data (as discussed above) and our synthetic data (SVReID) for training. The experiments are conducted with the baseline and the proposed method, respectively.
Impact of the training datasets. First, we compare the results using different real datasets (Mars, PRID, and LS-VID) for training. Specifically, on testing of iLIDS-VID, the largest LS-VID obtains the best performance. But we find that the performance of training on MARS is lower than that on PRID, although the videos and identifies in PRID are much less than the videos in MARS. However, the results on CCVID are significantly different from those on iLIDS-VID. We then analyze the cause of the above results. iLIDS-VID has various lighting and view-angle variations, which makes both MARS and PRID perform poorly. MARS shows worse generalization since its distribution is obviously different from that of iLIDS-VID, which has also been discussed in MARS paper (Zheng et al., 2016). The LS-VID with more various scenes and identities makes the model perform better. Different from iLIDS-VID, CCVID is built upon a gait recognition dataset captured by horizontal cameras with similar view angles, which is similar to MARS. Thus on CCVID, the training on MARS performs better than other datasets. The above results demonstrate that the performance for cross-domain ReID depends not only on the training data scale but also (more importantly) on its variety, especially the data style and distribution difference between the training and testing datasets. Even though Mars and LS-VID are large-scale, they can not cover the various scenarios during testing.
@@ 4.4 Cross-dataset evaluation results
Impact of the proposed method. We further compare the results of each row in Table 4.3. We can see that our method outperforms the baseline significantly under each setting of training and testing datasets. This clearly indicates that the proposed method can improve the cross-domain robustness not only for the scenes from synthetic to real domain but also those from real to real datasets.
Table 5: Experimental results of our method on the SVReID and the enlarged SVReID+ datasets. (%)
@@
4.5 Enlarged SVReID dataset
We then consider the data scale of the proposed SVReID. We can see that compared with some previous large datasets, such as Mars, and LS-VID, the scale of SVReID is not very large. As discussed before, SVReID has a promising generalization ability compared with other real video Re-ID datasets. Even so, due to the low cost of synthetic data generation, we further enlarge the proposed dataset to unleash the potential of synthetic data, in this section. Specifically, we enlarge it from 333 identities into 612, and from 1,164 videos into 2,823, to form SVReID+. As shown in Table 5, with training on SVReID+, the performance still improves as the data scale expands. However, we can see that, with the identities and videos enlarging approximately/over two times, the time of training will rise a lot, but the performance increase is not very significant. On the whole, considering the training time and the storage of training data, the SVReID is more economical and promising for the cross-domain video person Re-ID task.
4.6 Qualitative analysis
@@
5 Conclusion
In this work, we have studied a new and valuable problem of cross-domain video person Re-ID with synthetic data. For this problem, we developed a new framework, including a domain-invariant feature learning module and a mean-teacher network training scheme, both of which are achieved in a self-supervised manner without extra annotations. We have also built the benchmark including the training/evaluation datasets and comparative results, for this problem. In the future, we plan to build a unified and large synthetic dataset, including the clothes unaltering/changing, and other scenarios, for more general video person Re-ID training, for which we also consider using the recently popular generation models.
Data availability statements. The code and data in this work is publicly available at https://github.com/XiangqunZhang/UDA_Video_ReID.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under Grants: U1803264 and 62072334, the Hong Kong Innovation and Technology Fund under Grant: MHP/117/21.

### v3

@@ Methods
CAViT [15] 13.8 24.9 - 6.5 4.0 7.3 - 6.7
PTSA [16] 34.2 50.2 56.4 20.1 23.7 42.3 52.3 30.5
AP3D [10] 44.7 61.3 66.5 25.0 22.7 39.3 50.0 31.7
TCLNet [13] 42.6 57.7 64.9 22.8 26.7 50.0 61.3 34.9
CAL [12] 43.6 58.3 63.6 23.4 25.3 41.3 51.7 31.7
SINet [11] 42.8 57.6 63.1 23.4 29.3 47.3 63.3 39.8
GLT [21] 38.5 55.3 62.9 24.0 12.0 26.0 34.0 20.0
QAconv [25] 40.7 55.6 62.4 22.7 20.0 38.0 48.0 29.4
DCCT [19] 42.8 58.9 65.3 23.3 24.7 45.3 61.6 32.4
S2P [32] 32.4 48.9 55.2 19.7 18.7 35.3 43.4 27.0
TF-CLIP [20] 45.7 60.5 67.5 25.7 31.3 53.7 68.7 42.1
@@
CAViT [15] 13.5 38.2 - 25.2 12.2 20.5 - 4.1
PTSA [16] 47.9 71.4 80.1 51.8 10.6 18.1 22.5 3.7
AP3D [10] 56.2 80.9 92.1 67.4 13.4 21.7 26.1 4.5
TCLNet [13] 65.6 74.5 77.9 57.5 11.0 19.2 23.6 3.9
CAL [12] 51.6 73.0 78.7 55.1 11.4 20.3 25.2 4.0
SINet [11] 53.7 75.3 83.2 63.0 13.8 21.7 26.4 4.7
GLT [21] 33.7 62.9 71.9 48.0 9.1 16.7 20.9 3.1
QAconv [25] 50.6 77.5 84.3 61.8 11.3 20.6 25.6 4.1
DCCT [19] 49.7 74.7 81.4 59.5 14.3 22.5 28.3 5.3
S2P [32] 40.4 67.5 84.3 53.5 10.1 18.1 22.3 3.4
TF-CLIP [20] 63.4 72.4 88.7 67.8 16.2 25.6 31.8 5.7
@@
TABLE II: Comparison with the state-of-the-art Re-ID methods from the synthetic SVReID with clothes changing, i.e., SVReID (w cc), to the real dataset CCVID. (%)
@@
CAViT [15] 61.4 70.0 - 42.5
PTSA [16] 61.4 74.3 80.0 52.6
TCLNet [13] 42.8 57.6 63.1 23.4
CAL [12] 70.7 78.1 81.6 66.8
SINet [11] 70.7 79.1 83.1 64.1
GLT [21] 68.2 77.6 80.9 59.1
QAconv [25] 55.4 70.4 75.3 37.8
AP3D [10] 52.8 62.6 68.2 43.1
DCCT [19] 66.7 78.9 80.3 57.8
S2P [32] 53.4 65.7 72.3 41.3
TF-CLIP [20] 72.2 81.3 84.5 68.3
@@
TABLE III: Ablation study of the proposed method. (%)
@@
+ $\mathcal{L}^{\mathrm{S}}_{1}$ loss 32.7 59.3 73.3 45.4 71.7 78.3 81.8 68.1
+$\mathcal{L}^{\mathrm{S}}_{1}$ + $\mathcal{L}^{\mathrm{S}}_{2}$ losses 34.0 62.7 75.3 47.3 72.9 80.5 83.1 68.2
+ $\mathcal{L}^{\mathrm{S}}_{1}$ + $\mathcal{L}^{\mathrm{S}}_{2}$ + $\mathcal{L}^{\mathrm{S}}_{3}$ losses 36.0 63.3 75.3 49.1 73.9 79.9 84.2 70.1
w/o $\mathcal{L}^{\mathrm{C}}_{1}$ loss 39.0 62.7 73.3 50.1 75.2 82.4 86.8 71.8
w/o $\mathcal{L}^{\mathrm{C}}_{2}$ loss 39.3 64.0 78.0 51.2 74.6 80.3 83.8 70.0
@@
IV-B Comparison with state-of-the-art methods
Comparison methods. Due to limited research on cross-domain person Re-ID based on videos, we conducted a comprehensive evaluation by comparing our method against two categories of approaches. Similar with the division in the related work section, the first category is the state-of-the-art video-based person Re-ID approaches, including CAViT [15], PSTA [16], AP3D [10], TCLNet [13], SINet [11], CAL [12], DCCT [19] and TF-CLIP [20]. The second category is image-based person Re-ID methods that aim to handle the unsupervised domain adaptation (UDA) problem, including GLT [21], QAconv [25] and S2P [32].
As shown in Table I, we compare our method with the state-of-the-art video person Re-ID approaches and domain adaptive Re-ID approaches. We conduct the experiments on four pairs of cross-domain datasets. We can first see that the proposed method gets the best performance compared to other approaches with a large margin, on all four real-world datasets. Specifically, we find that our method outperforms the second-best approach (TF-CLIP) by over 10% in terms of mAP on iLIDS-VID and PRID datasets. The performance improvement on the large-scale Mars dataset also achieves over 5%. We can also see that on LS-VID the results of all methods are generally poor. This is because LS-VID is a large-scale dataset with more challenging testing cases.
As shown in Table II, on the CCVID dataset, the clothes-changing setting increases the difficulty for Re-ID. But since the data variety of CCVID (like the view angles, and illumination) is relatively simple, the overall performance on it is generally high. In this case, the proposed method still performs better than CAL, which is specifically designed for CCVID to handle the clothes-changing problem. These results demonstrate that the proposed method has an effective domain adaption ability across synthetic and real-world datasets.
IV-C Ablation study
To verify the effectiveness of each component in the proposed method, we conduct the ablation study on two datasets, i.e., the medium-scale dataset iLIDS-VID and the clothes-changing aware dataset CCVID, as shown in Table III.
Effectiveness of domain-invariant feature learning module (Section III-B). We add the self-supervised domain discrimination losses, i.e., $\mathcal{L}^{\mathrm{S}}_{1}$, $\mathcal{L}^{\mathrm{S}}_{2}$, and $\mathcal{L}^{\mathrm{S}}_{3}$ in Eq. (10) one by one. As shown in Table III, we can first see that the proposed self-supervised loss $\mathcal{L}^{\mathrm{S}}$ is effective in improving the Re-ID performance with a large margin. Also, the results demonstrate that each part in $\mathcal{L}^{\mathrm{S}}$ can improve the performance, respectively.
TABLE IV: Comparison results with the real-to-real data training. (%)
@@ IV-C Ablation study
Effectiveness of cross-domain consistency learning module (Section III-C). With $\mathcal{L}^{\mathrm{S}}$ equipped, we further investigate the effectiveness of the mean-teacher scheme. We can first see that the integration of the mean-teacher training strategy can further improve the performance of our method. Also, we remove each corresponding loss component i.e., $\mathcal{L}^{\mathrm{C}}_{1}$ in Eq (14) and $\mathcal{L}^{\mathrm{C}}_{2}$ in Eq. (15) in the self-supervised ID consistency learning, respectively. We can also see that each sub-loss in $\mathcal{L}^{\mathrm{C}}$ is verified to be useful.
@@
IV-D Cross-dataset evaluation results
In this section, we also use two real datasets, i.e., iLIDS-VID and CCVID for testing, and use the other three real datasets, i.e., Mars, PRID, and LS-VID for training. As shown in Table IV-C, we compare the cross-domain Re-ID results using the real data (as discussed above) and our synthetic data (SVReID) for training. The experiments are conducted with the baseline and the proposed method, respectively.
Impact of the training datasets. First, we compare the results using different real datasets (Mars, PRID, and LS-VID) for training. Specifically, on testing of iLIDS-VID, the largest LS-VID obtains the best performance. But we find that the performance of training on MARS is lower than that on PRID, although the videos and identifies in PRID are much less than the videos in MARS. However, the results on CCVID are significantly different from those on iLIDS-VID. We then analyze the cause of the above results. iLIDS-VID has various lighting and view-angle variations, which makes both MARS and PRID perform poorly. MARS shows worse generalization since its distribution is obviously different from that of iLIDS-VID, which has also been discussed in MARS paper [45]. The LS-VID with more various scenes and identities makes the model perform better. Different from iLIDS-VID, CCVID is built upon a gait recognition dataset captured by horizontal cameras with similar view angles, which is similar to MARS. Thus on CCVID, the training on MARS performs better than other datasets. The above results demonstrate that the performance for cross-domain ReID depends not only on the training data scale but also (more importantly) on its variety, especially the data style and distribution difference between the training and testing datasets. Even though Mars and LS-VID are large-scale, they can not cover the various scenarios during testing.
@@ IV-D Cross-dataset evaluation results
Impact of the proposed method. We further compare the results of each row in Table IV-C. We can see that our method outperforms the baseline significantly under each setting of training and testing datasets. This clearly indicates that the proposed method can improve the cross-domain robustness not only for the scenes from synthetic to real domain but also those from real to real datasets.
TABLE V: Experimental results of our method on the SVReID and the enlarged SVReID+ datasets. (%)
@@
IV-E Enlarged SVReID dataset
We then consider the data scale of the proposed SVReID. We can see that compared with some previous large datasets, such as Mars, and LS-VID, the scale of SVReID is not very large. As discussed before, SVReID has a promising generalization ability compared with other real video Re-ID datasets. Even so, due to the low cost of synthetic data generation, we further enlarge the proposed dataset to unleash the potential of synthetic data, in this section. Specifically, we enlarge it from 333 identities into 612, and from 1,164 videos into 2,823, to form SVReID+. As shown in Table V, with training on SVReID+, the performance still improves as the data scale expands. However, we can see that, with the identities and videos enlarging approximately/over two times, the time of training will rise a lot, but the performance increase is not very significant. On the whole, considering the training time and the storage of training data, the SVReID is more economical and promising for the cross-domain video person Re-ID task.
IV-F Qualitative analysis
@@
V Conclusion
In this work, we have studied a new and valuable problem of cross-domain video person Re-ID with synthetic data. For this problem, we developed a new framework, including a domain-invariant feature learning module and a mean-teacher network training scheme, both of which are achieved in a self-supervised manner without extra annotations. We have also built the benchmark including the training/evaluation datasets and comparative results, for this problem.
A limitation of this work seems that the used training and testing datasets for cloth-changing video person Re-ID is not large and various enough, although the cloth-changing setting is not the main purpose of this work. In the future, we plan to build a unified and large synthetic dataset, including the clothes unaltering/changing, and other scenarios, for more general video person Re-ID training, for which we also consider using the recently popular diffusion models.
This work was supported by the National Natural Science Foundation of China (NSFC) under Grant 62072334, 62402490 and the Shenzhen Basic Research Foundation under Grant JCYJ20240813154920027.